Ken is 3 times as old as Emma . Sam is half as old as Emma . The sum of their ages are 63. How old is each of them?

Answer by pj66(7)

Let x be Emma's age.
=> Ken's age = 3x and Sam's age = x/2
Also, 3x+x+(x/2)=63
=> 9x/2=63
=> 9x=126
=> x=14
Hence, Emma's age is 14 ,Ken's age = 42 and Sam's age is 7